# Architecting mechanisms of damage in topological metamaterials

### v2

@@ 5.2 Additional constraints
Figure 9: Effect of pinned boundary conditions. Maximum normalised stress in the elements cross section for (a) LP1 and (b) LUP1 at SR = 30. Damage path with pinned boundary for: (c) LP1 at SR = 10 and 30; (d) LP1 at SR = 30 with 10 unit pre-crack length; (e) LP2 at SR =30; (f) LUP1 at SR = 30; and, (g) LUP1 at SR = 30 with 50 units along the length. In (c-g) we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows either side of these zones. If the SSS can’t develop, we only plot the direction of the SSS with a single arrow. Maximum normalised bending stress in the elements cross section at SR = 30 for: (h) LP1 with 10 unit pre-crack length; (i) LUP1; and, (j) LUP1 with 50 units along the length.
@@
The lattice size should not significantly affect its dominant behaviour when the SSS at the long-wavelength limit dominate the mechanical response. This occurs when $l_{\mathrm{\theta}}$ is above a critical length that localises the SSS towards this limit, as these modes do not decay. In D.1 we simulate a series of LP1 lattices with increasing domain size but identical aspect ratio. Keeping the aspect ratio constant ensures the SSS zones relative size remains constant. Plotting the axial energy portion for these lattices shows that the dominant behaviour does not vary significantly, and therefore, for this aspect ratio the SSS associated with the long-wavelength limit dominates the response.
@@
A similar damage evolution is observed in other confined non-topologically polarised lattices when $z>2d$ (Hedvard et al., 2024). In these lattices, an incomplete debonding type failure occurs, leaving bridging elements along the interface. Subsequently, these elements fracture, completely debonding the interface. SSS are expected to form within the lattice as it is confined and $z>2d$. Therefore, this initial failure process may occur due to the removal of these SSS. Once removed, the remaining bridging elements become bending-dominated (i.e. the local $z\leq 2d$), leading to the observed toughening behaviour. For lattices with $z<2d$ this bridging is not as readily observed (Hedvard et al., 2024).
When the size of the lattice prevents the SSS from developing, the dominant behaviour and failure do not vary greatly between roller or pinned boundary conditions. This is the case for LP2, where the damage path is identical for both boundary conditions (Fig. 9e) and the axial energy only rises from $3\%$ to $11\%$. LP2 would require a minimum length of 49 units for the SSS at the long-wavelength limit to develop. Increasing the lattice length to 50 units results in these SSS developing, and in D.2 we show that the fracture initiates at the top and bottom edges. The axial energy portion further increases to $22\%$, highlighting that the original length of LP2 was not large enough for these SSS to develop.
For LUP1 the axial energy portion goes from approximately 70% to 95% when the boundary conditions are switched to pinned for SR = 30. The SSS long-wavelength limit zones shown in Fig. 9f strongly correlate to the stress that develops within the lattice (Fig. 9b). The SSS corresponding to $\theta_{1}$ and $\theta_{2}$ require 2 and 34 units along the length to develop for the domain size simulated. The SSS associated with $\theta_{2}$, therefore, does not occur. The change in the stress profile with respect to the form of these modes is provided in the D.3 for LUP1-3.
@@
6 Conclusions
Designing the mechanisms of damage in metamaterials has predominately leaned on the inclusion of heterogeneities to overcome the plethora of influencing parameters and competing length scales. In this work, we instead show that homogeneous Maxwell lattices have topological and geometry-dependent parameters that dominate the locations of stress concentrations, particularly at large SR. These parameters are derived from the form of the FM and SSS that exponentially localise on bulk discontinuities. The localisation of these modes is captured through a topological index, defining the equivalence class and constraining the microstructure to those which exhibit an identical topological protected directional preference. Within each equivalence class, the form of these modes i.e. the decay and angle of propagation, can be tuned by distorting the microstructure without affecting its topological polarisation $\textbf{P}_{\mathrm{T}}$. Using a beam element simulator, we numerically probed the effect of these modes over a range of slenderness for a set of distorted kagome lattices that span the possible equivalence classes.
@@ 6 Conclusions
When the lattice is bending-dominated, we show that the bending stress distribution within the lattice can be predicted through the bulk-boundary correspondence, which determines the amplitude of a disturbance in the bulk given a prescription on a boundary. For stretching-dominated lattices at low SR, while the stress profile within the bulk of the lattice is not governed by the FM that localise at the discontinuity, the bending stresses around the crack tip are. Consequently, when the fracture in these lattices is governed by bending, the damage strongly aligns with these modes. At low SR, the effect of these modes was found to diminish due to the delocalisation of the stress associated with increasing bending rigidity.
@@
Over a range of SR, we provide insight into the dominant behaviour and failure of topologically distinct kagome lattices. At large SR we are able to accurately map the stress and damage to the underlying microstructure, advancing our understanding of the mechanisms of damage in these lattice-based metamaterials. These findings provide a robust framework that can be leveraged to architect damage paths themselves. Moving towards a continuum description, this provides the necessary physics to inform the fracture process of these metamaterials.
Acknowledgments
Funding:
@@ Acknowledgments
Competing interests:
There are no competing interests to declare.
Appendix A Method
@@
A.1 Discretisation
Axial spring $i$ connect adjacent sites $j$ and $k$; and, torsional spring $s$ connects adjacent bonds $i$ and $l$ (Fig. 2a). When more than two bonds are connected at a site, torsional springs are added between all adjacent bonds (Fig. 2b). In all simulations, we discretise each bond into three elements.
A.2 Spring stiffness
@@
$k_{\mathrm{i}}=\frac{EA}{l_{\mathrm{i}}^{\mathrm{0}}}$ (9)
$c_{\mathrm{s}}=\frac{EI}{\min(l_{\mathrm{i}}^{\mathrm{0}},l_{\mathrm{l}}^{\mathrm{0}})}$ (10)
Where $E$ is the elastic modulus, $A$ is the cross sectional area, $I$ is the second moment of area and $l_{\mathrm{i/l}}^{\mathrm{0}}$ are the initial length of the elements. The length of bond $i$ and $l$, which are connected by torsional spring $s$ may not be equivalent. In this case, the minimum length of the two bonds was used to determine $c_{\mathrm{s}}$. Little difference was observed when taking the average length or maximum length.
A.3 Stress in bonds
@@ A.4 Algorithm
Appendix B Evanescent Mode Visualizations
In this section, we explain the methods used to visualize the form of evanescent modes associated with bulk discontinuities. It is important to mention that although the solutions we visualize given a surface disturbance are exact, the (functional) form of this disturbance has to be prescribed following certain assumptions. The visualizations are thus an instructive means of designing the lattices. For instance, in the case of FM on the crack surface of LP lattices, we use the qualitative assumption of a strongly localised disturbance on the crack tip, which can be motivated through fracture mechanics on micropolar media (Berkache et al., 2022), as well as our simulation results (see Fig.2i). For RP lattices, which do not possess a significant spike, we have instead used a sigmoid prescription, which aims to emulate the relaxation of the lattice on the crack-edge, and the fact that the disturbance approaches a uniform actuation in the bulk. Finally, although our examples are focused on FM visualisations, note that the same method can be applied for SSS.
B.1 Bulk-Boundary Correspondence
In the limit of linear elasticity, the existence of uniform strain, zero energy (e.g. GH) modes imply the existence of spatially modulated versions of the same strain that also cost no energy (Rocklin et al., 2017). For idealized, zero bending stiffness kagome lattices, the zero energy distortion on the unit cell can be captured by a single parameter $\phi$ (Sun et al., 2012; Nassar et al., 2020), which, when allowed to vary spatially defines a type of micro-deformation field $\phi(\boldsymbol{x})$. In the linear elastic limit, geometric compatibility has to be enforced, which constrains the functional form for $\phi(\boldsymbol{x})$, effectively reducing the degrees of freedom of the function from two independent coordinates $\phi\equiv\phi(x,x)$ into a combination of the two $\phi\equiv\phi(x+\lambda x)$.
@@ B.1 Bulk-Boundary Correspondence
allowing us to write, e.g. $q_{\mathrm{2}}\equiv G(q_{\mathrm{1}})$, prescribing one wavenumber component in terms of the other. Given that $C(q_{\mathrm{1}},q_{\mathrm{2}})$ contains all information about the lattice, this model does not make any homogenising assumptions - $q_{\mathrm{2}}=G(q_{\mathrm{1}})$ is exact, $\forall q_{\mathrm{1}}\in\left[-\pi/a,\pi/a\right]$. Note that, as $\mathbf{C}=\mathbf{Q}^{\dagger}$, the $q_{\mathrm{2}}$ component of the SSS will be the complex conjugate of the FM resulting in $\kappa^{\mathrm{SSS}}=-\kappa^{\mathrm{FM}}$ and $\theta^{\mathrm{SSS}}=\theta^{\mathrm{FM}}=\theta$—although we focus on FM, the extension to the discussion of SSS is straightforward. Now, using this result, consider the deformation field $\phi(\boldsymbol{x})$, written as an inverse Fourier Transform:
[…]
@@ B.3 Boundary Conditions
A FM, in general, distorts the lattice in a way that does not respect the Boundary Conditions (BC). An equivalent way of viewing the continuation of the disturbance around the boundary constraint is to consider a secondary disturbance $\phi_{b}(x,y)$ localized on the boundary, which, when superimposed with the primary disturbance localized on the crack tip $\phi_{0}(x,y)$, has to give a displacement field that respects the BCs. For the strongly localised crack tip actuation, there are two sets of Dirichlet BCs we use in this work: (1) rollers ($\delta y=0$) and (2) pinned ($\delta x=0$, $\delta y=0$). Although pinning the sites lifts the FM (Kane and Lubensky, 2014), the roller boundary conditions may allow for their existence by a reflection-like propagation that we examine below.
@@
For y-mirror-symmetric unit cells (LP, RP, TK), roller boundary conditions can be satisfied by superimposing two images of the disturbance, mirrored around the boundary. This is because the mirroring achieves the desired $\delta y_{b}=-\delta y_{0}$, on the boundary (see Fig. 10b). Furthermore, the continuation of the real domain into its virtual counterpart is compatible with the unit cell geometry, as the unit cell is chiral. For achiral unit cells that still belong in the equivalence classes (±,±,±)plus-or-minusplus-or-minusplus-or-minus($\pm$,$\pm$,$\pm$)( ± , ± , ± ) or (,±,±)minus-or-plusplus-or-minusplus-or-minus($\mp$,$\pm$,$\pm$)( ∓ , ± , ± ), mirroring will cause a discontinuity in the lattice tiling around the boundary. Nonetheless, mirroring would still satisfy the BC, with the extension of the FM from one domain to the other refracting the disturbance. Numerically, the FM can be constructed together with its reflection by obtaining the primary amplitude on the boundary $\phi_{0}(x,y_{b})$, and then using this as the new prescription for the secondary mode $\phi_{b}(x,y)$.
@@
For lattices in the (±,,±)plus-or-minusminus-or-plusplus-or-minus($\pm$,$\mp$,$\pm$)( ± , ∓ , ± ) or (±,±,)plus-or-minusplus-or-minusminus-or-plus($\pm$,$\pm$,$\mp$)( ± , ± , ∓ ) equivalence classes and initial geometries with a horizontal crack, like the ones considered in this work, both modes localize on the same horizontal edge, and the disturbance cannot be canceled on the boundary by a mode localizing there. It is not directly clear what combination of modes localizing on the remaining edges is induced in the lattice, so as to respect the boundary conditions. As a result, we limit our visualizations to the case of open boundary conditions (e.g. Fig. 10a) for these modes.
Appendix C Behaviour of additional topologically distinct lattices

### v3

@@ V.2 Additional constraints
Figure 9: Effect of pinned boundary conditions. Maximum normalised stress in the elements cross section for (a) LP1 and (b) LUP1 at SR = 30. Damage path with pinned boundary for: (c) LP1 at SR = 10 and 30; (d) LP1 at SR = 30 with 10 unit pre-crack length; (e) LP2 at SR =30; (f) LUP1 at SR = 30; and, (g) LUP1 at SR = 30 with 50 units along the length. In (c-g) we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows either side of these zones. If the SSS can’t develop, we only plot the direction of the SSS with a single arrow. Maximum normalised bending stress in the elements cross section at SR = 30 for: (h) LP1 with 10 unit pre-crack length; (i) LUP1; and, (j) LUP1 with 50 units along the length. The simulations for these lattices are included in movie S5.
@@
The lattice size should not significantly affect its dominant behaviour when the SSS at the long-wavelength limit dominate the mechanical response. This occurs when $l_{\mathrm{\theta}}$ is above a critical length that localises the SSS towards this limit, as these modes do not decay. In Appendix D.4 we simulate a series of LP1 lattices with increasing domain size but identical aspect ratio. Keeping the aspect ratio constant ensures the SSS zones relative size remains constant. Plotting the axial energy portion for these lattices shows that the dominant behaviour does not vary significantly, and therefore, for this aspect ratio the SSS associated with the long-wavelength limit dominates the response.
@@
A similar damage evolution is observed in other confined non-topologically polarised lattices when $z>2d$ [39]. In these lattices, an incomplete debonding type failure occurs, leaving bridging elements along the interface. Subsequently, these elements fracture, completely debonding the interface. SSS are expected to form within the lattice as it is confined and $z>2d$. Therefore, this initial failure process may occur due to the removal of these SSS. Once removed, the remaining bridging elements become bending-dominated (i.e. the local $z\leq 2d$), leading to the observed toughening behaviour. For lattices with $z<2d$ this bridging is not as readily observed [39].
When the size of the lattice prevents the SSS from developing, the dominant behaviour and failure do not vary greatly between roller or pinned boundary conditions. This is the case for LP2, where the damage path is identical for both boundary conditions (Fig. 9e) and the axial energy only rises from $3\%$ to $11\%$. LP2 would require a minimum length of 49 units for the SSS at the long-wavelength limit to develop. Increasing the lattice length to 50 units results in these SSS developing, and in Appendix F.1 we show that the fracture initiates at the top and bottom edges. The axial energy portion further increases to $22\%$, highlighting that the original length of LP2 was not large enough for these SSS to develop.
For LUP1 the axial energy portion goes from approximately 70% to 95% when the boundary conditions are switched to pinned for SR = 30. The SSS long-wavelength limit zones shown in Fig. 9f strongly correlate to the stress that develops within the lattice (Fig. 9b). The SSS corresponding to $\theta_{1}$ and $\theta_{2}$ require 2 and 34 units along the length to develop for the domain size simulated. The SSS associated with $\theta_{2}$, therefore, does not occur. The change in the stress profile with respect to the form of these modes is provided in Appenix F.2 for LUP1-3.
@@
VI Conclusions
Designing the mechanisms of damage in metamaterials has predominately leaned on the inclusion of heterogeneities to overcome the plethora of influencing parameters and competing length scales. In this work, we instead show that homogeneous Maxwell lattices have topological and geometry-dependent parameters that dominate the locations of stress concentrations, particularly at large SR. These parameters arise from the form of evanescent FM and SSS, which themselves derive from bulk geometrical properties, hence providing a design framework that does not rely on heterogeneities, and is agnostic with respect to the location of crack initiation. Furthermore, by finely controlling the form of these modes, i.e. their decay and angle of propagation, we demonstrate how our framework can be used to design the response of lattices without necessarily changing their topological properties. Using a beam element simulator, we numerically probed the effect of these modes over a range of slenderness for a set of distorted kagome lattices that span the possible equivalence classes.
@@ VI Conclusions
When the lattice is bending-dominated, we show that the bending stress distribution within the lattice can be predicted through the bulk-boundary correspondence, which determines the amplitude of a disturbance in the bulk given a prescription on a boundary. For stretching-dominated lattices at large SR, while the stress profile within the bulk of the lattice is not governed by the FM that localise at the discontinuity, the bending stresses around the crack tip are. Consequently, when the fracture in these lattices is governed by bending, the damage strongly aligns with these modes. At low SR, the effect of these modes was found to diminish due to the delocalisation of the stress associated with increasing bending rigidity.
@@
Over a range of SR, we provide insight into the dominant behaviour and failure of topologically distinct kagome lattices. At large SR we are able to accurately map the stress and damage to the underlying microstructure, advancing our understanding of the mechanisms of damage in these lattice-based metamaterials. Although only two boundary conditions are examined, the form of the edge states are independent of these, making the findings applicable to a wider range of boundary conditions. These findings provide a robust framework that can be leveraged to architect damage paths themselves. Moving towards a continuum description, this provides the necessary physics to inform the fracture process of these metamaterials.
Acknowledgements.
@@
Appendix A Nomenclature
Symbols
AA Cross sectional area of bonds
$A_{edge}$ Projected area of the edge $A_{edge}=bL$
$\mathbf{C}$ Compatibility matrix
$C_{\mathrm{s}}$ Torsional spring stiffness
$E$ Strain energy
$E_{\mathrm{m}}$ Elastic modulus
$F$ Applied force
$F_{\mathrm{tol}}$ Tolerance for convergence of FIRE 2.0 algorithm
$H$ Initial height of the lattice
II Second moment of area
$K_{\mathrm{i}}$ Axial spring stiffness
$L$ Length of the domain
NN Number of sites (nodes) in the lattice
$N_{\mathrm{c}}$ Number of constraints in the lattice
$N_{\mathrm{0}}$ Number of floppy modes (FM)
$N_{\mathrm{s}}$ Number of states of self-stress (SSS)
$P_{\mathrm{T}}$ Topological polarisation vector
$\mathbf{Q}$ Equilibrium matrix
SR Slenderness ratio, SR $|a_{1}|/t$
$\textbf{a}_{1}$, $\textbf{a}_{2}$, $\textbf{a}_{3}$ Primitive lattice vectors
$b$ Depth of bonds
$\mathbf{d}$ Dimensions
$\mathbf{e}$ Bond elongation vector
$\mathbf{f}$ Site force vector
$l_{\mathrm{i}}$, $l_{\mathrm{i}}^{\mathrm{0}}$ Current and original lengths of axial springs
$l_{\mathrm{\theta}}$ Length of the boundary within which the SSS at the long wavelength limit may develop
$\textbf{p}_{1}$, $\textbf{p}_{2}$, $\textbf{p}_{3}$ Vectors normal to the bonds
$q_{\mathrm{1}},q_{\mathrm{2}}$ Wavenumber components in the 1 and 2 directions
$\textbf{r}_{1}^{\mathrm{0}}$, $\textbf{r}_{2}^{\mathrm{0}}$, $\textbf{r}_{3}^{\mathrm{0}}$ Basis vectors for the sites of the RK
$\mathbf{t}$ Bond tension vector
$t$ Thickness of bonds
$\mathbf{u}$ Site displacement vector
$x_{\mathrm{\mu}}$ Distortion parameter affecting the $\mathrm{\mu}$-th filament
$y_{\mathrm{j}}$ Site degrees of freedom
$z$ Coordination number of a lattice
$z_{\mathrm{s}}$ Triangle scaling parameter in kagome lattice
$\alpha_{\mathrm{GH}}$ Expected angle between bonds $i$ and $l$ under deformation following the GH mode
$\alpha_{\mathrm{s}}$, $\alpha_{\mathrm{0}}^{\mathrm{s}}$ Current and original angles in torsional springs
$\delta$ Applied displacement
$\epsilon_{\mathrm{n}}$ Nominal strain, $\epsilon_{\mathrm{n}}=\delta/H$
$\kappa_{\mathrm{1}},\kappa_{\mathrm{2}}$ Imaginary part of the wave numbers in the 1 and 2 directions
$\phi$ Amplitude of the floppy mode in the bulk
$\sigma_{\mathrm{A}}$, $\sigma_{\mathrm{B}}$ Axial and bending stresses in lattice elements
$\tilde{\sigma}_{\mathrm{A}}$, $\tilde{\sigma}_{\mathrm{B}}$ Normalised axial and bending stresses in lattice elements, $\tilde{\sigma}_{\mathrm{A}}$, $\tilde{\sigma}_{\mathrm{B}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}$, $\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}$
$\sigma_{\mathrm{A}}^{\mathrm{u}}$, $\sigma_{\mathrm{B}}^{\mathrm{u}}$ Material failure stresses in axial and bending
$\sigma_{n}$ Normalized applied stress
$\theta$ Propagation angle of floppy modes
Abbreviations
AR Aspect Ratio
D-FM Domain wall of Floppy Modes
D-SSS Domain wall of States of Self-Stress
DW Domain wall
FM Floppy Modes
GH Mode Guest-Hutchinson Mode
LP Left Polarised
LPU Left-Up Polarised
LTK Left Twisted kagome
NN Nearest Neighbour
NNN Next Nearest Neighbour
RK Regular kagome lattice
RP Right Polarised
RDP Right-Down Polarised
RTK Right Twisted kagome
SSS States of Self-Stress
TK Twisted kagome
Appendix B Lattice definitions
A summary of the lattices simulated within this study are provided in Fig. 10.
Figure 10: Lattice definition with values of the parametrisation $\{x_{\mathrm{1}},x_{\mathrm{2}},x_{\mathrm{3}}\}$ for: (a) Regular kagome (RK); (b) Twisted kagome (TK); (c) Right polarised (RP); (d) Left polarised (LP); (e) Right-down polarised (RDP); and, Left-up polarised (LP)
Appendix C Method
@@
C.1 Discretisation
Axial spring $i$ connect adjacent sites $j$ and $k$; and, torsional spring $s$ connects adjacent bonds $i$ and $l$ (Fig. 2a). When more than two bonds are connected at a site, torsional springs are added between all adjacent bonds (Fig. 2b). In all simulations, we discretise each bond into three elements. This level of discretization cannot resolve modes beyond the second buckling mode. However, given the quasi-static loading conditions and the constraints applied in our simulations, higher order deformation modes are not expected to occur.
C.2 Spring stiffness
@@
$K_{\mathrm{i}}=\frac{E_{\mathrm{m}}A}{l_{\mathrm{i}}^{\mathrm{0}}}$ (9)
$C_{\mathrm{s}}=\frac{E_{\mathrm{m}}I}{\min(l_{\mathrm{i}}^{\mathrm{0}},l_{\mathrm{l}}^{\mathrm{0}})}$ (10)
Where $E_{\mathrm{m}}$ is the elastic modulus, $A$ is the cross sectional area, $I$ is the second moment of area and $l_{\mathrm{i/l}}^{\mathrm{0}}$ are the initial length of the elements. The length of bond $i$ and $l$, which are connected by torsional spring $s$ may not be equivalent. In this case, the minimum length of the two bonds was used to determine $C_{\mathrm{s}}$. Little difference was observed when taking the average length or maximum length.
C.3 Stress in bonds
@@ C.4 Algorithm
Appendix D Evanescent Mode Visualizations
In this section, we explain the methods used to visualize the form of evanescent modes associated with bulk discontinuities. It is important to mention that although the solutions we visualize given a surface disturbance are exact, the (functional) form of this disturbance has to be prescribed following certain assumptions. The visualizations are thus an instructive means of designing the lattices. For instance, in the case of FM on the crack surface of LP lattices, we use the qualitative assumption of a strongly localised disturbance on the crack tip, which can be motivated through fracture mechanics on micropolar media [6], as well as our simulation results (see Fig.2i). For RP lattices, which do not possess a significant spike, we have instead used a sigmoid prescription, which aims to emulate the relaxation of the lattice on the crack-edge, and the fact that the disturbance approaches a uniform actuation in the bulk. Finally, although our examples are focused on FM visualisations, note that the same method can be applied for SSS.
D.1 Bulk-Boundary Correspondence
In the limit of linear elasticity, the existence of uniform strain, zero energy (e.g. GH) modes imply the existence of spatially modulated versions of the same strain that also cost no energy [38]. For idealized, zero bending stiffness kagome lattices, the zero energy distortion on the unit cell can be captured by a single parameter $\phi$ [40, 41], which, when allowed to vary spatially defines a type of micro-deformation field $\phi(\bm{x})$. In the linear elastic limit, geometric compatibility has to be enforced, which constrains the functional form for $\phi(\bm{x})$, effectively reducing the degrees of freedom of the function from two independent coordinates $\phi\equiv\phi(x,y)$ into a combination of the two $\phi\equiv\phi(x+\lambda y)$.
@@ D.1 Bulk-Boundary Correspondence
allowing us to write, e.g. $q_{\mathrm{2}}\equiv G(q_{\mathrm{1}})$, prescribing one wavenumber component in terms of the other. Given that $C(q_{\mathrm{1}},q_{\mathrm{2}})$ contains all information about the lattice, this model does not make any homogenising assumptions - $q_{\mathrm{2}}=G(q_{\mathrm{1}})$ is exact, $\forall q_{\mathrm{1}}\in\left[-\pi/a,\pi/a\right]$. Note that, as $\mathbf{C}=\mathbf{Q}^{\dagger}$, the $q_{\mathrm{2}}$ component of the SSS will be the complex conjugate of the FM resulting in $\kappa^{\mathrm{SSS}}=-\kappa^{\mathrm{FM}}$ and $\theta^{\mathrm{SSS}}=\theta^{\mathrm{FM}}=\theta$—although we focus on FM, the extension to the discussion of SSS is straightforward. Now, using this result, consider the deformation field $\phi(\bm{x})$, written as an inverse Fourier Transform:
$\phi(x,y)$ $=\int_{-\infty}^{\infty}dq_{\mathrm{1}}e^{iq_{\mathrm{1}}x}\int_{-\infty}^{\infty}dq_{\mathrm{2}}e^{iq_{\mathrm{2}}y}\ \tilde{\phi}(q_{\mathrm{1}},q_{\mathrm{2}})\delta(q_{\mathrm{2}}-G(q_{\mathrm{1}}))$
$=\int_{-\infty}^{\infty}dq_{\mathrm{1}}\tilde{\phi}(q_{\mathrm{1}})e^{iq_{\mathrm{1}}x}e^{iG(q_{\mathrm{1}})y},$ (14)
@@ D.3 Boundary Conditions
A FM, in general, distorts the lattice in a way that does not respect the Boundary Conditions (BC). An equivalent way of viewing the continuation of the disturbance around the boundary constraint is to consider a secondary disturbance $\phi_{b}(x,y)$ localized on the boundary, which, when superimposed with the primary disturbance localized on the crack tip $\phi_{0}(x,y)$, has to give a displacement field that respects the BCs. For the strongly localised crack tip actuation, there are two sets of Dirichlet BCs we use in this work: (1) rollers ($\delta y=0$) and (2) pinned ($\delta x=0$, $\delta y=0$). Although pinning the sites lifts the FM [20], the roller boundary conditions may allow for their existence by a reflection-like propagation that we examine below.
@@
For y-mirror-symmetric unit cells (LP, RP, TK), roller boundary conditions can be satisfied by superimposing two images of the disturbance, mirrored around the boundary. This is because the mirroring achieves the desired $\delta y_{b}=-\delta y_{0}$, on the boundary (see Fig. 11b). Furthermore, the continuation of the real domain into its virtual counterpart is compatible with the unit cell geometry, as the unit cell is chiral. For achiral unit cells that still belong in the equivalence classes (±,±,±)($\pm$$\pm,\pm$) or $(\mp,\pm,\pm)$, mirroring will cause a discontinuity in the lattice tiling around the boundary. Nonetheless, mirroring would still satisfy the BC, with the extension of the FM from one domain to the other refracting the disturbance. Numerically, the FM can be constructed together with its reflection by obtaining the primary amplitude on the boundary $\phi_{0}(x,y_{b})$, and then using this as the new prescription for the secondary mode $\phi_{b}(x,y)$.
Figure 12: Dominant behaviour and failure of additional lattices. Snapshot of: (a) axial energy portion in the entire lattice; and, (b) axial stress fraction in the element that fails preceding the first failure event. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for RTK at: (c) SR = 3; and, (d) 30. The difference in angle is visualised through the colour directly to the right of each unit cell. Damage path for: (e) RTK; and, (f) RDP1. The damage process is visualised with a line traversing the domain. The line connects separated nodes, where fracture occured, within the lattice. For clarity, we do not show the micro-structure throughout the domain. The damage process for SR = 3, 10, 30 is plotted with green, black and blue lines and associated rectangular markers respectively (refer to (b)). Arrows indicate the direction of damage. Normalised maximum bending stress in cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (g, k) RDP1 (0.1,-0.1,0.1); and, (h, l) RDP3 (0.16,-0.04,0.16) at SR = 30. (i) Damage path for RDP1 and RDP3 at SR = 30. (j) $\theta$ spectrum for RDP1 and RDP3.
@@
For lattices in the $(\pm,\mp,\pm)$ or $(\pm,\pm,\mp)$ equivalence classes and initial geometries with a horizontal crack, like the ones considered in this work, both modes localize on the same horizontal edge, and the disturbance cannot be canceled on the boundary by a mode localizing there. It is not directly clear what combination of modes localizing on the remaining edges is induced in the lattice, so as to respect the boundary conditions. As a result, we limit our visualizations to the case of open boundary conditions (e.g. Fig. 11a) for these modes.
D.4 Dominant SSS
Appendix E Behaviour of additional topologically distinct lattices